steam at 100c is passed into 1.1 kg of water contained in a calorimeter of water equivalent 0.02kg at 15c till the temperature of calorimeter and its contents rises to 80c the mass of steam condensed is ?

steam at  100c is passed into 1.1 kg of water contained in a calorimeter of water equivalent 0.02kg at 15c till the temperature of calorimeter and its contents rises to 80c the mass of steam condensed is ?

Grade:

3 Answers

Adwait Mastud
7 years ago
student-name Priyanka asked in PhysicsChapter-Heat and Thermodynamics1.Steam at 100oC is passed into 1.1kg of water contained in a calorimeter of water equivalent 0.02kg at 15oC till the temperature of the calorimeter rises to 800C . The mass of steam condensed in kilogram is :(a)0.13 (b)0.065 (c)0.260 (d)0.135 0 Follow 0student-name Ruhhi Ralhan answered thisin Physics, ClassDear Student,Given:Mass of water and calorimeter=(1.10+0.02)=1.12 kgspecific heat capacity of water=4.184×103 Jkg-1K-1Now, using Q=mcΔTheat gained by calorimeter and water,Let the mass of steam=m kgLatent heat of vaporization of water at 100°C=540 Cal/g=540×4.184×103 Jkg-1Heat lost by steam,now equate above two equations
Ankit
6 years ago
Let the mass of sream = x When it is passed through water it release heat and that heat is gained by water to increase its temp.. So, heat released =heat gained by water x×540 +x×1×(100-80) = 1.1×1×(80-15) + 0.02 ×1×(80-15) On solving this we get x= 0.13 kg
Yash Chourasiya
3 years ago
Dear Student

We Know That
Heat lost by steam =Heat gained by (water +calorimeter)
mL + m × c × (100−80) = 1.12 × c × (80−15)
m[540 + 1 × 20] = 1.12 × 1 × 65
m = 0.13kg

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya